Suppose the multivariable demand for Good X was given by the function: QDx = 73 -Px -0.5 Py.

Where, Px represents the price of the good and Py represents the price of another good.

If the price of Good Y (Py) is $40, then what is the choke price of Good X?

(Do not include a $ sign in your response. Round to the nearest 2 decimal places if necessary.)
Answer #1

QDx = 73 - Px - 0.5 PY

if PY = 40, Then, QDx = 73 - Px - (0.5 40) = 53 - Px

Now, choke price is when QDx = 0 (No demand for the good)

Therefore the equation becomes 0 = 53 - Px

i.e., Px = 53

The choke price is 53. Below that price there is a demand for good x and above there is no demand al all.